Can someone please help me with this problem?
The population P in 1993 for a state given along with r, its annual percentage rate of continuous growth
P= 34 millions, r= 2.6%
(a) Write the formula f(x) = P e^yx, where r is in the decimal notation, that models the population in millions x years after 1991.
(b) Estimate the population in 2010
F(x) = ?
The population in 2015 will be approximately how many million?
(round to the nearest tenth as needed)
Thank you so much for your time. I appreciate it Answer by nerdybill(7384) 
 The population P in 1993 for a state given along with r, its annual percentage rate of continuous growth
P= 34 millions, r= 2.6%
(a) Write the formula f(x) = P e^yx, where r is in the decimal notation, that models the population in millions x years after 1991.
.
.
Don't know whether you meant 1993 or 1991 (I'll assume 1991)
.
Given:
f(x) = P e^(yx)
where
f(x) is popuplation x years after 1991
y is .026
P is 34 million
.
your equation is then:
f(x) = 34e^(.026x)
.
(b) Estimate the population in 2010
F(x) = ?
x = 2010 - 1991
x = 19
f(19) = 34e^(.026*19)
f(19) = 34*1.63886
f(19) = 55.7 million
.
The population in 2015 will be approximately how many million?
(round to the nearest tenth as needed)
.
x = 2015 - 1991
x = 24
f(24) = 34e^(.026*24)
f(24) = 34*1.86638
f(24) = 63.5 million